If I add 65 grams of calcium chloride to 500 grams of water, what will the boiling



points be of the resulting solution? Kb = 0.512 °C.

1
Expert's answer
2022-04-26T17:14:02-0400

CaCl2 —> Ca2+ + 2Cl-


ΔTb= i. Kb⋅b, where

i - the van't Hoff factor;

Kb - the ebullioscopic constant of the solvent;

b - the molality of the solution. 

ΔTb - the poiling point elevation - defined as Tb − Tb°

The ebullioscopic constant of water is equal to 0.512C kg mol-1


To determine the molality of the solution, use calcium chloride's molar mass to help you find the number of moles you have in that 65-g sample.


n(CaCl2) = m / Mr = 65 g / 110.98 g/mol = 0.59 moles


This means that the molality of the solution will be - do not forget to convert the mass of the solvent from grams to kilograms

b = nsolute / msolvent

b = 0.59 moles / 0.500 kg = 1.18 molal

i (CaCl2) = 3

ΔTb= 3 •0.512C kg mol-1• 1.18 molal= 1.81°C

ΔTb = Tb - Tb°

Tb = 100 °C + 1.81°C = 101.81°C